# Finiteness of leaps of modules of integrable derivations of algebras of finite type

### v1

@@
Abstract
We prove the finiteness of leaps of modules of integrable derivations for algebras essentially of finite type and, more generally, for schemes essentially of finite type over an algebraically closed field of positive characteristic. This gives an affirmative answer to a question posed by L.N. Macarro.
1 Introduction
Hasse-Schmidt derivations (abbreviated as HS-derivations) are introduced in [SchmidtHasse+1937+215+237], which are also called higher derivations [Mat, §27]. Given a positive integer $m$ and a $k$-algebra $R$, $\operatorname{HS}_{k}^{m}(R)$ denotes the group of HS-derivations of length $m$, which consists of $A_{m}$-algebra automorphisms
$R\otimes_{k}A_{m}\to R\otimes_{k}A_{m},\ A_{m}:=k[t]/(t^{m+1})$
that becomes the identity when restricted to $R\otimes_{k}(A_{m}/(t))=R$. For $m=\infty$, we set $A_{\infty}=k[\![t]\!]$ and we also have the notion of HS-derivations of length $\infty$. There is a canonical isomorphism between $\operatorname{HS}_{k}^{1}(R)$ and the module $\operatorname{Der}_{k}(R,R)$. Under this identification, the image of
@@ 1 Introduction
is called the module of $m$-integrable derivations and denoted by $\operatorname{Der}_{k}^{m}(R)$, which has been studied in [matsumura1982integrable], [NARVAEZMACARRO20122712], [NARVAEZMACARRO2024109758]. Note that if $k$ contains $\mathbb{Q}$ or if $R$ is formally smooth over $k$, we have $\operatorname{Der}_{k}^{m}(R)=\operatorname{Der}_{k}(R)$ for any $m$ (see [Mat]). Thus, the study of these modules is expected to lead to a better understanding of singularities in positive characteristic. We have the following descending chain
@@
We would like to know when the inclusion $\operatorname{Der}_{k}^{m-1}(R)\supset\operatorname{Der}_{k}^{m}(R)$ is proper. When this is the case, we say that $R$ leaps at $m$. It has been known that leaps occur only at $m=p^{i}$ where $p$ is the characteristic of $k$ ([Her, Theorem 4.1]). The number of leaps is finite if $R$ is a unibranch curve ([NARVAEZMACARRO2024109441, Theorem 4.7]), and, more generally, if $R$ is a reduced curve ([bravo2024finitenessleapssensehasseschmidt, Theorem 6.1, Corollary 6.3]). The aim of this paper is to show that, using results from [miy2025], the finiteness of leaps holds for every algebra essentially of finite type over an algebraically closed field of positive characteristic. Precisely, we show:
@@
$\operatorname{Der}_{k}^{M}(R)=\operatorname{Der}_{k}^{M+1}(R)=\dots=\operatorname{Der}_{k}^{\infty}(R).$
This theorem provides, for the first time, a general finiteness result for leaps in arbitrary dimension, thereby giving an affirmative answer to Macarro’s question [8, Question 3.6.5] and contributing to a better structural understanding of HS-derivations. Since leaps are stable under separable base changes by [hernandez2020behavior, Theorem 3.27], this also implies the finiteness of leaps over any perfect field. As an application of Theorem 1.1, we also give an affirmative answer to another question [NARVAEZMACARRO20122712, Question 3.6.1] by the following:
@@
Let $R$ be a $k$-algebra essentially of finite type, $n\geq 1$ be an integer and $D\in\operatorname{HS}_{k}^{n}(R)$. Then $D$ is $\infty$-integrable if and only if $D$ is $m$-integrable for every integer $m\geq n$. Here, we say that $D$ is $m$-integrable if $D$ is in the image of $\operatorname{HS}_{k}^{m}(R)\to\operatorname{HS}_{k}^{n}(R)$.
In Section 2, we recall the definition of HS-derivations and results that are going to be used later.
In Section 3, we introduce the notion of $a$-boundedness of HS-derivations. Intuitively, this concept represents how each term of a HS-derivation is shrinking with respect to the $\mathfrak{m}$-adic topology.
In Section 4, we prove the finiteness of leaps in the local case. The idea is that, starting from a bounded HS-derivation, we construct another one that is less bounded and coincides with the original one in small degrees.
In Section 5, we prove the finiteness of leaps in the global case and discuss its consequences.
@@
Acknowledgements. The author would like to thank Professor Keiji Oguiso and the members of Oguiso’s laboratory for their interest in this work and their valuable comments and suggestions.
2 Preliminaries
We first recall basic facts of HS-derivations (cf. [SchmidtHasse+1937+215+237], [NARVAEZMACARRO20122712, Chapter 1], [Mat, §27]) and results on obstruction modules for HS-derivations (cf. [NARVAEZMACARRO20122712], [Her], [miy2025, §2]).
Definition 2.1 (= [NARVAEZMACARRO20122712, Definition 1.2.1-1.2.6]).
Let $R$ be a $k$-algebra. Let $m\geq 1$ be an integer or $m=\infty$. Then a Hasse-Schmidt derivation of length $m$ is a sequence $D=(D_{i})_{i=0}^{m}$ of $k$-linear endomorphisms $R\to R$ such that $D_{0}=\operatorname{id}$ and that
@@ 2 Preliminaries
for every $x,y\in R,i\geq 1.$ Let $E=(E_{i})_{i=0}^{m}$ be another Hasse-Schmidt derivation. We define the composition of $D$ and $E$ by
@@
The symbol $\operatorname{HS}_{k}^{m}(R)(=\operatorname{HS}_{k}(R,m))$ denotes the set of all the Hasse-Schmidt derivations of length $m.$ It is a group whose multiplication is given by $\circ$. Let $x\in R$. We define
@@
A Hasse-Schmidt derivation $D=(D_{i})_{i=0}^{m}$ defines a natural $A_{m}$-algebra automorphism
$R\otimes_{k}A_{m}\to R\otimes_{k}A_{m},\ x\otimes 1\mapsto\sum_{i=0}^{m}D_{i}(x)\otimes t^{i}$
where $A_{m}:=k[t]/t^{m+1}$ for $m<\infty$ and $A_{\infty}:=k[\![t]\!]$. This gives an isomorphism of groups between $\operatorname{HS}_{k}^{m}(R)$ and the group of $A_{m}$-algebra automorphisms of $R\otimes_{k}A_{m}$ that becomes the identity when restricted to $R\otimes_{k}(A_{m}/(t))=R$. Let $1\leq l\leq m$. We define the truncation map
@@ A Hasse-Schmidt derivation $D=(D_{i})_{i=0}^{m}$ defines a natural $A_{m}$-algebra automorphism
In this paper, we usually denote the image again by D, by abuse of notation. It is a group homomorphism and commutes with $x\bullet$ for $x\in R$. Note that the set $\operatorname{HS}_{k}^{1}(R)$ can be identified with the $R$-module $\operatorname{Der}_{k}(R,R)$ under
@@
Definition 2.2 (= [NARVAEZMACARRO20122712, Definition 2.1.1], [matsumura1982integrable, §1]).
@@
It is a sub-$R$-module of $\operatorname{Der}_{k}(R,R)$.
Let $d\in\operatorname{Der}_{k}(R,R)$. For $1<m<\infty$, we say that $d$ leaps at $m$ if $d\in\operatorname{Der}_{k}^{m-1}(R,R)$ and $d\notin\operatorname{Der}_{k}^{m}(R,R)$. If $d\in\cap_{1<m<\infty}\operatorname{Der}_{k}^{m}(R,R)$, then $d$ does not leap. Let $S\subset\operatorname{Der}_{k}(R,R)$ be a subset. By $\operatorname{Leaps}(S)\subset\mathbb{N}$, we denote the set of leaps produced by $S$. The set of leaps of $R$ is $\operatorname{Leaps}(\operatorname{Der}_{k}(R,R))$.
Theorem 2.3 (= [NARVAEZMACARRO20122712, Corollary 2.3.5]).
@@
Thus, for a $k$-scheme $X$ essentially of finite type, its $m$-integrable derivations consist a sheaf denoted by $\operatorname{Der}_{k}^{m}(\mathcal{O}_{X})$ for $m<\infty$.
Theorem 2.4 (= [Her, Theorem 4.1]).
@@
Let us recall the first cotangent module $T^{1}_{R/k}$ of a $k$-algebra $R$ essentially of finite type (cf. [Ser, Definition 1.1.6]). Let $T$ be a localization of a polynomial algebra over $k$ and $I\subset T$ an ideal such that $T/I=R.$ By [Ser, Corollary 1.1.8], we have the following canonical exact sequence of $R$-modules
@@
Proposition 2.6 (= [miy2025, Proposition 2.12]).
In the situation of Definition 2.5, the map $\operatorname{ob}_{m}$ is a group homomorphism that does not depend on the choice of $T,I,E$. A Hasse-Schmidt derivation $D\in\operatorname{HS}_{k}^{m-1}(R)$ extends to an element in $\operatorname{HS}_{k}^{m}(R)$ if and only if $\operatorname{ob}_{m}(D)=0$. For every $D\in\operatorname{HS}_{k}^{m-1}(R)$ and $x\in R$, we have
@@
Definition 2.7 (= [miy2025, Definition 2.13]).
Let $R$ be a $k$-algebra essentially of finite type and $m\geq 2$ an integer. We define the $m$-th obstruction module $\operatorname{Ob}^{m}_{R}\subset T^{1}_{R/k}$ as the image of $\operatorname{ob}_{m}$. At this point, this is an abelian group.
Proposition 2.8 (= [miy2025, Proposition 2.14]).
@@ 3 On $a$-boundedness
for every $n,n^{\prime}$]. When $\mathfrak{m}$ is clear from the context (e.g., when $R$ is a local ring and $\mathfrak{m}$ is the maximal ideal), we simply say that $D$ is $a$-bounded [resp. logarithmically $a$-bounded].
Note that the composition of logarithmically $a$-bounded HS-derivations is again logarithmically $a$-bounded.
@@
First, note that $\widetilde{a}$ is also subadditive. We show that by induction of $n^{\prime}=0,1,\dots,$ we have
@@ 4 The local case
$(T,\mathfrak{m})$ is a formally smooth local $k$-algebra that is a localization of a polynomial ring in finite variables at a prime ideal (not necessarily maximal) and $I\subset\mathfrak{m}$ is an ideal of $T$. Set $(R,\mathfrak{n})=(T/I,\mathfrak{m}/I)$; this is a local $k$-algebra essentially of finite type. $X=\operatorname{Spec}R$ and $U=X\setminus\{\mathfrak{n}\}$ are $k$-schemes essentially of finite type and $\iota:U\hookrightarrow X$ is an open immersion. The $k$-scheme $U$ satisfies $\operatorname{FL}(\tau)$ for an integer $\tau\geq 1$. We write $q=p^{\tau}$.
Note that we include the case where $X$ consists of one point, where $U=\emptyset$ satisfies $\operatorname{FL}(\tau)$ for obvious reasons. The goal of this section is to show that $X$ itself satisfies $\operatorname{FL}(\theta)$ for another sufficiently large integer $\theta$.
@@
Let $M:=I/I^{2}$ and $H_{n}:=\operatorname{Hom}_{R}(M,\mathfrak{n}^{n})$ for $n=1,2,\dots.$ Consider $H_{n}$ as submodules of $H:=\operatorname{Hom}_{R}(M,R)$. Then there exists a positive integer $C_{1}$ such that $H_{n}\subset\mathfrak{n}^{n-C_{1}}\cdot H$ for $n\geq C_{1}$.
@@
Let $R^{e_{1}}\to R^{e_{2}}\to M\to 0$ be a finite presentation for $M$. For each $n=1,2,\dots$, we have the following commutative diagram
@@
From this diagram, we see that $H_{n}=H\cap(\mathfrak{n}^{n})^{e_{2}}$ where the intersection is taken in $R^{e_{2}}$. We apply the Artin-Rees lemma to $H\subset R^{e_{2}}$ with respect to the $\mathfrak{n}$-adic topology and the result follows. ∎
@@
Let $M:=I/I^{2}$ and $G$ be the image of the natural homomorphism $\operatorname{Hom}_{R}(\Omega_{T/k}\otimes_{T}R,R)\to\operatorname{Hom}_{R}(M,R)=H$. Then there exists a positive integer $C_{2}$ such that $G\cap(\mathfrak{n}^{n}\cdot H)\subset\mathfrak{n}^{n-C_{2}}\cdot G$ for $n\geq C_{2}$.
@@
This is also a consequence of the Artin-Rees lemma. ∎
@@
Let $i\geq 0$ be an integer, $S$ be a $k$-algebra and $J\subset S$ an ideal. Define
$J^{[p^{i}]}:=\{x^{p^{i}}|x\in J\}$
which is an ideal of $S^{p^{i}}$. It is the image of $J$ under the surjection $S\to S^{p^{i}}$. In particular, we have a natural inclusion $J^{[p^{i}]}S\subset J^{p^{i}}$ of ideals of $S.$ Also, observe that $(J^{j})^{[p^{i}]}=(J^{[p^{i}]})^{j}$ for any $j\geq 0$.
@@
There exists a positive integer $C_{3}$ such that, for every integer $i\geq\tau$, we have $\operatorname{Ob}_{R}^{p^{i}}\cap(\mathfrak{n}^{nq}\cdot T_{R/k}^{1})\subset(\mathfrak{n}^{n-C_{3}})^{[q]}\cdot\operatorname{Ob}_{R}^{q}$ if $n\geq C_{3}$.
@@
by 2.11. Hence, it suffices to show the existence of $C_{3}$ such that $F\cap(\mathfrak{n}^{nq}\cdot T_{X/k}^{1})\subset(\mathfrak{n}^{n-C_{3}})^{[q]}\cdot\operatorname{Ob}_{X}^{q}$ for $n\geq C_{3}$. Take $c$ as above so that $(\mathfrak{n}^{[q]})^{n-c}R\supset\mathfrak{n}^{nq}$ for every $n\geq c$. By the Artin-Rees lemma, there exists $c_{1}\geq 0$ such that
@@
if $n\geq c_{1}$. As the quotient sheaf $F/\operatorname{Ob}_{X}^{q}$ is equal to zero over $U$, there exists $c_{2}\geq 0$ such that
@@
There exists a positive integer $C_{4}$ such that, for every $i=0,1,\dots,\tau-1$, we have $\operatorname{Ob}_{R}^{p^{i}}\cap(\mathfrak{n}^{np^{i}}\cdot T_{R/k}^{1})\subset(\mathfrak{n}^{n-C_{4}})^{[p^{i}]}\cdot\operatorname{Ob}_{R}^{p^{i}}$ if $n\geq C_{4}$. Here, we set $\operatorname{Ob}_{R}^{1}=0$.
@@
It suffices to show the claim when $i$ is fixed. Take $c_{3}$ so that $(\mathfrak{n}^{[p^{i}]})^{n-c_{3}}R\supset\mathfrak{n}^{p^{i}n}$ for every $n\geq c_{3}$. By the Artin-Rees lemma, there exists $c_{4}\geq 0$ such that

### v4

@@
Abstract.
We prove the finiteness of leaps of modules of $m$-integrable derivations for algebras essentially of finite type and, more generally, for schemes essentially of finite type over an algebraically closed field of positive characteristic. This provides an affirmative answer to a question posed by L. Narváez Macarro. As an application, we establish the coherence of the module of $\infty$-integrable derivations.
2020 Mathematics Subject Classification:
13N15, 13A35, 14F10
1. Introduction
Hasse-Schmidt derivations (abbreviated as HS-derivations) were introduced in [13], which are also called higher derivations [6, §27]. Given a positive integer $m$, a field $k$ and a $k$-algebra $R$, $\operatorname{HS}_{k}^{m}(R)$ denotes the group of HS-derivations of length $m$, which consists of $A_{m}:=k[\![t]\!]/(t^{m+1})$-algebra automorphisms
$R[\![t]\!]/(t^{m+1})\to R[\![t]\!]/(t^{m+1})$
that become the identity when restricted to $R[\![t]\!]/(t^{m+1})\otimes_{A_{m}}k=R$. For $m=\infty$, we regard $t^{\infty}=0$, and we also have the notion of HS-derivations of infinite length. There is a canonical isomorphism between $\operatorname{HS}_{k}^{1}(R)$ and the module $\operatorname{Der}_{k}(R,R)$. Under this identification, the image of
@@ 1. Introduction
is called the module of $m$-integrable derivations and denoted by $\operatorname{Der}_{k}^{m}(R)$, which has been studied in [5], [11], [12]. The module $\operatorname{Der}_{k}^{\infty}(R)$ is also denoted by $\operatorname{Ider}_{k}(R)$. Note that if $k$ contains $\mathbb{Q}$ or if $R$ is formally smooth over $k$, we have $\operatorname{Der}_{k}^{m}(R)=\operatorname{Der}_{k}(R)$ for any $m\leq\infty$ (see [6]). Thus, the study of integrable derivations is expected to lead to a better understanding of singularities in positive characteristic. From now on, we assume that $k$ is an algebraically closed field of positive characteristic $p$ and $R$ is essentially of finite type. We have the following descending chain
@@
The question of when the inclusion $\operatorname{Der}_{k}^{m-1}(R)\supset\operatorname{Der}_{k}^{m}(R)$ is proper has been studied since their introduction. When this is the case, we say that $R$ leaps at $m$. It has been known that leaps occur only at $m=p^{i}$ ([2, Theorem 4.1]). The number of leaps is finite if $R$ is a unibranch curve ([10, Theorem 4.7]), and, more generally, if $R$ is a reduced curve ([1, Theorem 6.1, Corollary 6.3]). The aim of this paper is to show, using results from [7, §2], that the finiteness of leaps holds for every algebra essentially of finite type over an algebraically closed field of positive characteristic. Precisely, we show:
@@
$\operatorname{Der}_{k}^{M}(R)=\operatorname{Der}_{k}^{M+1}(R)=\dots=\operatorname{Der}_{k}^{\infty}(R)=\operatorname{Ider}_{k}(R).$
This theorem provides, for the first time, a general finiteness result for leaps in arbitrary dimensions, thereby giving an affirmative answer to a question posed by L. Narváez Macarro [8, Question 3.6.5] and contributing to a better structural understanding of HS-derivations. Since leaps are stable under separable base changes by [3, Theorem 3.27], this also implies the finiteness of leaps over any perfect field. We also give an affirmative answer to another question [11, Question 3.6.1] by the following:
@@
Let $n\geq 1$ be an integer and let $D\in\operatorname{HS}_{k}^{n}(R)$. Then $D$ is $\infty$-integrable if and only if $D$ is $m$-integrable for every integer $m\geq n$. Here, we say that $D$ is $m$-integrable if $D$ is in the image of $\operatorname{HS}_{k}^{m}(R)\to\operatorname{HS}_{k}^{n}(R)$.
In fact, Theorem 1.1 is a special case of Theorem 1.2 where we put $n=1$. As an application of Theorem 1.1, we obtain:
Corollary 1.3 (=Corollary 5.6).
The module $\operatorname{Der}_{k}^{\infty}(R)=\operatorname{Ider}_{k}(R)$ is compatible with localization of $R$. That is, we have
$\operatorname{Ider}_{k}(S^{-1}R)=S^{-1}\operatorname{Ider}_{k}(R).$
This is the first coherence result for $\operatorname{Ider}_{k}(R)$ to the best of our knowledge. For $m<\infty$, a similar statement has been proven in [11, Corollary 2.3.5]. The case $m=\infty$ is more subtle because the map in [11, Corollary 1.3.6] is not surjective for $m=\infty$ while $m$ is assumed to be an integer there (cf. [15, Example 1.4]). Note that Corollary 1.3 implies that $\operatorname{Ider}_{k}(R)$ constitutes a coherent sheaf $\operatorname{Ider}_{k}(\mathcal{O}_{X})$ on an algebraic scheme $X$ over $k$. In particular, it is expected to behave better than the usual tangent sheaf. In fact, some pathological phenomena of derivations peculiar to positive characteristic do not occur if we replace it by integrable derivations. For example, the local rank of $\operatorname{Ider}_{k}(R)$ is at most $\operatorname{dim}R$, while it is not true for $\operatorname{Der}_{k}(R)$ (see [9]). This pathology does not occur in characteristic zero (see [4] when $R$ is a domain and [8] in general). We can also state Theorem 1.1 in terms of schemes:
Theorem 1.4 (= Theorem 5.7).
Let $X$ be a $k$-scheme essentially of finite type (=locally essentially of finite type and quasi-compact). Then there exists $M>0$ such that
$\operatorname{Der}_{k}^{M}(\mathcal{O}_{X})=\operatorname{Der}_{k}^{M+1}(\mathcal{O}_{X})=\dots=\operatorname{Ider}_{k}(\mathcal{O}_{X}),$
where $\operatorname{Der}_{k}^{m}(\mathcal{O}_{X})$ denotes the sheaf of $m$-integrable derivations (see Theorem 2.3).
We briefly outline the contents of the paper.
In Section 2, we recall the definition of HS-derivations and results that will be used later. A key relation is that the obstructions to extending HS-derivations are canonically dual to $m$-integrable derivations (Theorem 2.11). We also define the property $\operatorname{FL}(i)$ for algebras and schemes, which states that leaps do not occur at every $m>p^{i}$ (Definition 2.12).
In Section 3, we introduce the notion of $a$-boundedness for HS-derivations. Intuitively, this concept measures how each term of a HS-derivation shrinks with respect to the $\mathfrak{m}$-adic topology.
In Section 4, we prove the finiteness of leaps in the local case. In this section we fix a local $k$-algebra $(R,\mathfrak{n})$ such that $\operatorname{Spec}R\setminus\{\mathfrak{n}\}$ satisfies $\operatorname{FL}(\tau)$ for $\tau>0$, and we will prove that $\operatorname{Spec}R$ itself satisfies the finiteness of leaps. The idea is that, starting from a bounded HS-derivation, we modify it step by step so that the obstruction at each stage becomes zero while the lower terms remain unchanged. We use the Artin-Rees lemma to obtain positive constants which control the bounds. The essential step is Theorem 4.15, which states that elements of
$\mathfrak{n}^{e(q+1)+1}\cdot\operatorname{Der}_{k}^{q}(R)$
do not leap, where $q=p^{\tau}$ and $e\gg 0$. Then the fact that $R/\mathfrak{n}^{e(q+1)+1}$ is of finite length implies the finiteness of leaps of $R$ (Theorems 4.16, 4.17).
In Section 5, we prove the finiteness of leaps in the global case and discuss its consequences. We use noetherian induction to reduce to the case of Section 4 (Theorem 5.1). The finiteness of leaps of $m$-integrable derivations implies that of obstruction modules. This enables us to extend HS-derivations step by step (Lemma 5.3), and taking the limit, we obtain a HS-derivation of infinite length: This method yields a characterization of $\infty$-integrability (Theorem 5.4).
@@
2. Preliminaries
We first recall basic facts of HS-derivations (cf. [13], [11, Chapter 1], [6, §27]) together with results on $m$-integrable derivations and obstruction modules (cf. [11], [2], [7, §2]). The obstruction modules govern the extension of HS-derivations from length m1m-1 to $m$. These obstructions form coherent sheaves on schemes essentially of finite type (Theorem 2.10). Moreover, the ascending filtration of obstruction modules is naturally dual to the descending chain of $m$-integrable derivations (Theorem 2.11), a relation that plays a key role in establishing the finiteness of leaps.
Definition 2.1 (= [11, Definition 1.2.1-1.2.6]).
Let $R$ be a $k$-algebra. Let $m\geq 1$ be an integer or $m=\infty$. Then a Hasse-Schmidt derivation (abbreviated as HS-derivation) of length $m$ is a sequence $D=(D_{i})_{i=0}^{m}$ of $k$-linear endomorphisms $R\to R$ such that $D_{0}=\operatorname{id}$ and that
@@ 2. Preliminaries
for every $x,y\in R$, $i\geq 1.$ Let $E=(E_{i})_{i=0}^{m}$ be another HS-derivation. We define the composition of $D$ and $E$ by
@@
The symbol $\operatorname{HS}_{k}^{m}(R)(=\operatorname{HS}_{k}(R,m))$ denotes the set of all the HS-derivations of length $m.$ It forms a group under this compsition. Let $x\in R$. We define
@@
(Although $D[n]$ is usually defined to have length $mn$, it naturally extends to length mn+n1mn+n-1 by adding zero terms.) A HS-derivation $D=(D_{i})_{i=0}^{m}$ defines a natural $A_{m}$-algebra automorphism
$R[\![t]\!]/(t^{m+1})\to R[\![t]\!]/(t^{m+1}),\ x\mapsto\sum_{i=0}^{m}D_{i}(x)t^{i}\quad\mathrm{for}\ x\in R$
where $A_{m}:=k[\![t]\!]/t^{m+1}$. (We regard $t^{\infty}=0$.) This yields an isomorphism of groups between $\operatorname{HS}_{k}^{m}(R)$ and the group of $A_{m}$-algebra automorphisms of $R[\![t]\!]/(t^{m+1})$ that become the identity when restricted to $R[\![t]\!]/(t^{m+1})\otimes_{A_{m}}k=R$. Let $1\leq l\leq m$. We define the truncation map
@@
It is a group homomorphism and commutes with $x\bullet$ for $x\in R$. Note that for $m=1$, $\operatorname{HS}_{k}^{1}(R)$ identifies with the $R$-module $\operatorname{Der}_{k}(R,R)$ under
@@
Definition 2.2 (= [11, Definition 2.1.1], [5, §1]).
@@
It is a sub-$R$-module of $\operatorname{Der}_{k}(R,R)$. We also denote $\operatorname{Der}_{k}^{\infty}(R)$ by $\operatorname{Ider}_{k}(R)$.
Let $d\in\operatorname{Der}_{k}(R,R)$. For $1<m<\infty$, we say that $d$ leaps at $m$ if $d\in\operatorname{Der}_{k}^{m-1}(R)$ and $d\notin\operatorname{Der}_{k}^{m}(R)$. If $d\in\cap_{1<m<\infty}\operatorname{Der}_{k}^{m}(R)$, then $d$ does not leap. Let $S\subset\operatorname{Der}_{k}(R,R)$ be a subset. By $\operatorname{Leaps}(S)\subset\mathbb{N}$, we denote the set of leaps produced by $S$. The set of leaps of $R$ is $\operatorname{Leaps}(\operatorname{Der}_{k}(R,R))$.
Theorem 2.3 (= [11, Corollary 2.3.5]).
@@
This implies that if $m<\infty$ and $X$ is a $k$-scheme essentially of finite type, its $m$-integrable derivations consist a sheaf on $X$, which is denoted by $\operatorname{Der}_{k}^{m}(\mathcal{O}_{X})$.
Theorem 2.4 (= [2, Theorem 4.1]).
@@
Let us recall the first cotangent module $T^{1}_{R/k}$ of a $k$-algebra $R$ essentially of finite type (cf. [14, Definition 1.1.6]). Let $T$ be a localization of a polynomial algebra over $k$ and $I\subset T$ an ideal such that $T/I=R.$ By [14, Corollary 1.1.8], we have the following canonical exact sequence of $R$-modules
@@
We note that $\operatorname{ob}_{mn}(D[n])=\operatorname{ob}_{m}(D)$ for another integer $n>1$ because $E_{m}=E[n]_{mn}$ in the above construction.
Proposition 2.6 (= [7, Proposition 2.12]).
In the situation of Definition 2.5, the map $\operatorname{ob}_{m}$ is a group homomorphism independent of the choices of $T,I,E$. A HS-derivation $D\in\operatorname{HS}_{k}^{m-1}(R)$ extends to an element of $\operatorname{HS}_{k}^{m}(R)$ if and only if $\operatorname{ob}_{m}(D)=0$. For every $D\in\operatorname{HS}_{k}^{m-1}(R)$ and $x\in R$, we have
@@
Definition 2.7 (= [7, Definition 2.13]).
Let $R$ be a $k$-algebra essentially of finite type and $m\geq 2$ an integer. We define the $m$-th obstruction module $\operatorname{Ob}^{m}_{R}\subset T^{1}_{R/k}$ as the image of $\operatorname{ob}_{m}$. We set $\operatorname{Ob}_{R}^{1}=0\subset T^{1}_{R/k}$. At this point, this is an abelian group.
Proposition 2.8 (= [7, Proposition 2.14]).
@@ 3. On $a$-boundedness
for every $n,n^{\prime}$]. When $\mathfrak{m}$ is clear from the context (e.g., when $R$ is a local ring and $\mathfrak{m}$ is the maximal ideal), we simply say $a$-bounded [resp. logarithmically $a$-bounded].
Note that the composition of logarithmically $a$-bounded HS-derivations is again logarithmically $a$-bounded because $a$ is subadditive.
@@
First observe that $\widetilde{a}$ is also subadditive. We show that by induction of $n^{\prime}=0,1,\dots,$ we have
@@ 4. The local case
$(T,\mathfrak{m})$ is a formally smooth local $k$-algebra that is a localization of a polynomial ring in finite variables at a prime ideal (not necessarily maximal) and $I\subset\mathfrak{m}$ is an ideal of $T$. Set $(R,\mathfrak{n})=(T/I,\mathfrak{m}/I)$; this is a local $k$-algebra essentially of finite type. $X=\operatorname{Spec}R$ and $U=X\setminus\{\mathfrak{n}\}$ are $k$-schemes essentially of finite type and $\iota:U\hookrightarrow X$ is an open immersion. The $k$-scheme $U$ satisfies $\operatorname{FL}(\tau)$ for an integer $\tau\geq 1$. We put $q=p^{\tau}$.
Note that this also includes the case where $X$ consists of a single point, where $U=\emptyset$ satisfies $\operatorname{FL}(\tau)$ for obvious reasons. The aim of this section is to show that $X$ itself satisfies $\operatorname{FL}(\theta)$ for another sufficiently large integer $\theta$. The idea is to start with a bounded HS-derivation and modify it step by step so that the obstruction at each stage becomes zero while the lower terms remain unchanged. To achieve this, we first obtain certain explicit positive constants and use them to construct suitable subadditive sequences.
@@
Let $M:=I/I^{2}$ and let $H_{n}:=\operatorname{Hom}_{R}(M,\mathfrak{n}^{n})$ for $n=1,2,\dots.$ We regard $H_{n}$ as submodules of $H:=\operatorname{Hom}_{R}(M,R)$. Then there exists a positive integer $C_{1}$ such that
$H_{n}\subset\mathfrak{n}^{n-C_{1}}\cdot H\quad(\forall n\geq C_{1}).$
@@
Let $R^{e_{1}}\to R^{e_{2}}\to M\to 0$ be a finite presentation of $M$. For each $n=1,2,\dots$, we have the following commutative diagram
@@
From this diagram, we obtain $H_{n}=H\cap(\mathfrak{n}^{n})^{e_{2}}$ where the intersection is taken inside $R^{e_{2}}$. Applying the Artin-Rees lemma to the inclusion $H\subset R^{e_{2}}$ with respect to the $\mathfrak{n}$-adic filtration yields the claim. ∎
@@
Let $M:=I/I^{2}$ and let $G$ be the image of the natural homomorphism $\operatorname{Hom}_{R}(\Omega_{T/k}\otimes_{T}R,R)\to\operatorname{Hom}_{R}(M,R)=H$. Then there exists a positive integer $C_{2}$ such that
$G\cap(\mathfrak{n}^{n}\cdot H)\subset\mathfrak{n}^{n-C_{2}}\cdot G\quad(\forall n\geq C_{2}).$
@@
This also follows from the Artin-Rees lemma. ∎
@@
Let $i\geq 0$ be an integer, let $S$ be a $k$-algebra and let $J\subset S$ be an ideal. Define
$J^{[p^{i}]}:=\{x^{p^{i}}|x\in J\},$
an ideal of $S^{p^{i}}$. It is the image of $J$ under the surjection $S\to S^{p^{i}}$. In particular, we have a natural inclusion $J^{[p^{i}]}S\subset J^{p^{i}}$ of ideals of $S.$ Also, observe that $(J^{j})^{[p^{i}]}=(J^{[p^{i}]})^{j}$ for any $j\geq 0$.
@@
There exists a positive integer $C_{3}$ such that, for every integer $i\geq\tau$, we have
$\operatorname{Ob}_{R}^{p^{i}}\cap(\mathfrak{n}^{nq}\cdot T_{R/k}^{1})\subset(\mathfrak{n}^{n-C_{3}})^{[q]}\cdot\operatorname{Ob}_{R}^{q}\quad(\forall n\geq C_{3}).$
@@
by 2.11. Hence, it suffices to show the existence of $C_{3}$ such that $F\cap(\mathfrak{n}^{nq}\cdot T_{X/k}^{1})\subset(\mathfrak{n}^{n-C_{3}})^{[q]}\cdot\operatorname{Ob}_{X}^{q}$ for $n\geq C_{3}$. Choose $c$ as above so that $(\mathfrak{n}^{[q]})^{n-c}R\supset\mathfrak{n}^{nq}$ for every $n\geq c$. By the Artin-Rees lemma, there exists $c_{1}\geq 0$ such that
@@
if $n\geq c_{1}$. As the quotient sheaf $F/\operatorname{Ob}_{X}^{q}$ is supported at $\{\mathfrak{n}\}$, there exists $c_{2}\geq 0$ such that
@@
There exists a positive integer $C_{4}$ such that, for every $i=0,1,\dots,\tau-1$, we have
$\operatorname{Ob}_{R}^{p^{i}}\cap(\mathfrak{n}^{np^{i}}\cdot T_{R/k}^{1})\subset(\mathfrak{n}^{n-C_{4}})^{[p^{i}]}\cdot\operatorname{Ob}_{R}^{p^{i}}\quad(\forall n\geq C_{4}).$
@@
It suffices to show the claim when $i$ is fixed. Take $c_{3}$ so that $(\mathfrak{n}^{[p^{i}]})^{n-c_{3}}R\supset\mathfrak{n}^{p^{i}n}$ for every $n\geq c_{3}$. Applying the Artin-Rees lemma yields $c_{4}\geq 0$ such that